OpenStudy (anonymous):

( w+8) + (-4) rewrite using associative

OpenStudy (dayakar):

(w+8)+(-4) = W+ [8+(-4)]

OpenStudy (anonymous):

Can you simplify?

OpenStudy (dayakar):

(a+b)+c =a+(b+c)

OpenStudy (dayakar):

w+(8-4) =w+4

OpenStudy (anonymous):

I'm still confused on on rewritten part?

OpenStudy (dayakar):

where do u confused

OpenStudy (dayakar):

r u there

OpenStudy (denisaboichuk):

(w+8)+(-4) = W+ [8+(-4)]